OpenStudy (anonymous):

Please Solve for x y=x^(2)+2x-3

OpenStudy (anonymous):

try to factor it

OpenStudy (anonymous):

*and set it to 0

OpenStudy (anonymous):

So (x-3)(x+1)=y

OpenStudy (anonymous):

you can either factor or use quadratic equation whichever you prefer. factoring is faster so... -1 and 3 (x-1)(x-3)

OpenStudy (anonymous):

so x is 1 and 3

OpenStudy (anonymous):

but im trying to take the derivative of the inverse function

OpenStudy (anonymous):

Prob. should have said that in the beginning :)

OpenStudy (anonymous):

what is the inverse function of the equation?

OpenStudy (anonymous):

my homework says to use the formula for the derivative of the inverse function to compute f(-3)

OpenStudy (ranga):

To compute f(-3) or f^(-1) (-3) ? That is. f(-3) or evaluate inverse function at x = -3

OpenStudy (anonymous):

is the (2)+2x-3 an exponent the x?

OpenStudy (anonymous):

write it exactly how your hw says it

OpenStudy (anonymous):

You don't have to factor it if you know the 2 Tips for quick solving quadratic equations; Tip 1. When a + b + c = 0, one real root is (1) and the other is (c/a). In your example: x^2 + 2x - 3 = 0, a + b + c = 0, then one root is (1) and the other is (-3). TiP 2. When a - b + c = o, then one root is (-1) and the other is (-c/a).